Three resistors of resistance 5Ω, 3Ω and 4Ω are arranged in parallel between two terminals. Calculate the equivalent resistance.
1
Expert's answer
2020-06-22T11:11:05-0400

Equivalent resistance of N resistors arranged in parallel:


"R_T=1/\displaystyle\sum_{i=1}^N1/R_i"


"R_T = 1/(1/5 + 1/3 + 1/4) = 60/(12 + 20 + 15) = 60/47 \approx 1.277\quad(\Omega)"